• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name danez

    Log entry time 12:14:18 on March 07, 2011

    Entry number 350809

    keyword=Right Bogies Controls Acting Up

    This morning while attempting to move the right spectrometer, the bogies control for that arm wouldn't open properly. (The left bogies control operatated normally.)

    Upon resetting the right bogies IOC, the control GUI was able to be opened and used.

    While attempting to move the right spectrometer again in the afternoon, the same thing happened.